OpenStudy (anonymous):

Complete the square to find the standard form of the given equation. Then give the center and radius of the circle. X^2+y^2+16x-10y+8=0

OpenStudy (math&ing001):

This is how you complete the square : (x²+2*8x+8²) - 8² + (y²-2*5y+5²) - 5² = 0 Tell me if you need further help.

OpenStudy (anonymous):

is that the final answer?

OpenStudy (math&ing001):

No, it's not. You need to write it in form (x-a)²+(y-b)²=r² with (a,b) being the center and r the radius of the circle.

OpenStudy (anonymous):

can you please show me how to do that

OpenStudy (anonymous):

I am completely lost when it comes to these sort of problems

OpenStudy (math&ing001):

Ok, do you know this rule about quadratics a²+2ab+b²=(a+b)² ?

OpenStudy (anonymous):

yes

OpenStudy (math&ing001):

Can you apply it for (x²+2*8x+8²) and (y²-2*5y+5²) ?

OpenStudy (anonymous):

If you could give me the answer to the problem I will go over how to do it with you as soon as I submit the answer, it is due at 9:30 please help!!

OpenStudy (math&ing001):

We'll get it done before 9:30 I can promise you that.

OpenStudy (anonymous):

okay thank you

OpenStudy (math&ing001):

(x²+2*8x+8²)=(x+8)² and (y²-2*5y+5²) = (y-5)² The standard form will be (x+8)²+(y-5)²=89 Hence (-8,5) the center and √89 the radius.

OpenStudy (anonymous):

the answer says that everything is right except for the diameter and radius

OpenStudy (math&ing001):

Did you put the square root before 89 ?

OpenStudy (math&ing001):

√89=9.43

OpenStudy (anonymous):

Im on webwork so you type it in like sqrt89

OpenStudy (anonymous):

but if the diameter is wrong the radius will be too

OpenStudy (math&ing001):

Why did you get the diameter wrong ?

OpenStudy (anonymous):

I didnt. They are saying that 89 is the wrong answer

OpenStudy (nikato):

should be 81

OpenStudy (nikato):

\[x ^{2}+16x+y ^{2}^{-10y=-8}\]

OpenStudy (anonymous):

81 is right thank you

OpenStudy (nikato):

complete the square \[(x ^{2}+16+64)+(y ^{2}-10y+25)=-8+64+25\]

OpenStudy (math&ing001):

Of course it is I forgot the +8 -.-

OpenStudy (nikato):

simplify and factor \[(x+8)^{2}+(y-5)^{2}=81\]

OpenStudy (anonymous):

what would you radius be of a diameter with endpoints (4,5) (12,7) and a center (8,6)?

OpenStudy (anonymous):

the equation i have so far in standard form is (x-8)^2 + (y-6)^2 =?

OpenStudy (math&ing001):

Use the Distance Formula : d=sqrt( (x2-x1)² + (y2-y1)² ) between the points (12,7) and (8,6). 'd' will be the radius.

OpenStudy (imstuck):

Heres another way that might help. Begin by grouping the x and y terms together inside parenthesis, like this:\[(x ^{2}+16x)+(y ^{2}-10y)=-8\]and moving the 8 over to the other side.

OpenStudy (imstuck):

Now we can complete the square on the x terms and the y terms by taking half the x coefficient and squaring it and adding it inside the x parenthesis, and then adding it to the -8. Do the same with the y terms, to get this:\[(x ^{2}+16x+64)+(y ^{2}-10y+25)=-8+64+25\]

OpenStudy (imstuck):

Now you can make your perfect square binomials:\[(x+8)^{2}+(y-5)^{2}=81\]This looks like a circle to me with a center of (-8,5) and a radius of 9.